OpenStudy (anonymous):

For each function, y varies directly with x. Find each constant of variation. Then find the value of y when x= -0.3. 14. Y= 7 when x=2

OpenStudy (bibby):

y=kx 7=2k k=\(\dfrac{7}{2}\) \(y=kx, k=\dfrac{7}{2},x=-0.3\\y=\dfrac{7}{2}*-0.3=-1.05\)

OpenStudy (anonymous):

I'm so confused ^^^ ☹

OpenStudy (bibby):

when y varies directly with x, it means there's some constant (a number) k that when multiplied to x gives us y y=kx we know that Y= 7 when x=2 we plug that in and solve for k

OpenStudy (anonymous):

So I do 7= k (2) ?

OpenStudy (dayakar):



OpenStudy (anonymous):

Wait is the answer 3.5?

OpenStudy (dayakar):

y2= 7(-0.3)/2 = -2.1/2

OpenStudy (dayakar):

y2= -1.05

OpenStudy (anonymous):

What 😳^^

OpenStudy (dayakar):

\[y \alpha x\]

OpenStudy (dayakar):

y= kx y/x =k

OpenStudy (anonymous):

I have no clue what that means

OpenStudy (anonymous):

Can you check my other question and see if my answers are correct

OpenStudy (dayakar):

y directly varies x means if y increase x increases and vice versa is it clear

OpenStudy (anonymous):

So the answer to this problem is y2 = -1.05

OpenStudy (dayakar):



OpenStudy (dayakar):

get it

OpenStudy (anonymous):

I can't see that